what is the vapor pressure of a solution made by dissolving 225 grams of glucose in 358 mL of water at 30°C? The molar mass of glucose in 180.2 g/mol. What is the vapor pressure lowering? The vapor pressure of pure water at 30°C is 31.82 mmHg. Assume that density of the solution is 1.00 g/mL.


1
Expert's answer
2021-05-06T07:40:15-0400

Q189725


what is the vapor pressure of a solution made by dissolving 225 grams of glucose in 358 mL of water at 30°C? The molar mass of glucose is 180.2 g/mol. What is the vapor pressure lowering? The vapor pressure of pure water at 30°C is 31.82 mmHg. Assume that the density of the solution is 1.00 g/mL.


Solution :


Vapor pressure lowering is related to the mole fraction of solute in the solution by the formula.


"\Delta{P} _{solvent} = \chi_{solute} * P_{solvent}"


χ solute is the mole fraction of solute.


ΔPsolvent is the vapor pressure lowering.


Psolvent is the vapor pressure of the solvent.


First find the moles of glucose present in 225 grams.


"moles \ of glucose = 225 \cancel{\ g \ of \ glucose } * \frac{1 \ mol \ of \ glucose }{180.2 \cancel{ \ g \ of \ glucose } }"

"= 1.2486 \ mol \ of \ glucose"


Next, find the moles of water in 358 mL. The density of water is 1.00g/mL.


mass of water = density * volume = 1.00 g/mL * 358 mL


= 358 grams.



"moles \ of H_2O= 358 \cancel{\ g \ of \ H_2O } * \frac{1 \ mol \ of \ H_2O }{18.015 \cancel{ \ g \ of \ H_2O } }"


"= 19.872 \ moles \ of \ H_2O"



"moles \ fraction \ of \ glucose, \ \chi_{glucose} = \frac{moles \ of \ glucose }{moles \ of \ glucose \ + \ moles \ of \ H_2O}"


"\ \chi_{glucose} = \frac{1.2486 \ moles \ of \ glucose }{1.2486 \ moles \ of \ glucose \ + \ 19.872 \ moles \ of \ H_2O}"


"\ \chi_{glucose} = \frac{1.2486 }{21.1209} = 0.05912 ;"

Now we know,

χ glucose is = 0.05912 , and P water = 31.82 mm Hg.


glucose is the solute and water is the solvent here.



"\Delta{P} _{water } = \chi_{glucose } * P_{water}"


"\Delta{P} _{water } = 0.05912 \ * \ 31.82 \ mm \ Hg."

"\Delta{P} _{water } = 1.881 \ mm \ Hg ;"

In the correct significant figure, the answer is 1.88 mm Hg.


Hence the vapor pressure of water is lowered by 1.88 mm Hg.